OpenStudy (anonymous):

which equation is solved for the variable y? z=4x+2y y=z-2x y=z/2-2x y=z/2-4x y=z/2+2x

OpenStudy (anonymous):

y = z/2 - 4x

OpenStudy (anonymous):

no, its the second one

OpenStudy (anonymous):

z=4x+2y move 4x to the other side z-4x=2y divide both sides by 2 z/2-2x=y which is the second answer choice

OpenStudy (anonymous):

submitting....

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

got a C thx